How much work is required to accelerate a 1500kg car from 15 m/s to 40 m/s?What is the average net force acting on the car if it covers a distance of 275m? 


1
Expert's answer
2022-01-10T09:09:11-0500

Given:

"v_1=\rm 15\: m/s"

"v_2=\rm 40\: m/s"

"m=\rm 1500\: kg"

"d=\rm 275\: m"


The energy-work theorem says

"\frac{mv_2^2}{2}-\frac{mv_1^2}{2}=W=Fd"

(a) the work done

"W=\frac{mv_2^2}{2}-\frac{mv_1^2}{2}\\=\frac{1500}{2}(40^2-15^2)=1\:031\:250\: \rm J"

(b) the average force

"F=\frac{W}{d}=\frac{1031250}{275}=3\:750\: \rm N"